# Interpreting the Weight Space of Customized Diffusion Models

### v2

@@ 5 Limitations
As with any data-driven method, w2w space inherits the biases of the data used to discover it. For instance, co-occurring attributes in the identity-encoding models would cause linear classifier directions to entangle them (e.g. gender and facial hair). However, as we scale the number of models, spurious correlations will drop as evidenced by Fig. 10. These directions are also limited by the labels present in CelebA. Additionally, the span of the w2w space is dictated by the models used to create it. Thus, w2w space can struggle to represent more complex identities as seen in Fig. 11. Inversion in these cases amounts to projecting onto the closest identity on the weights manifold. Despite these limitations, our analysis on the size of the model dataset reveals that forming a space using a larger and more diverse set of identity-encoding models can mitigate this limitation.
@@ Appendix A Sampling
We present additional examples of models sampled from w2w space in Fig. 12. The sampled models encode a diverse array of identities which are not copied from the dataset of model weights, as seen by comparing them to the nearest neighbor models. However, there are attributes borrowed from the nearest neighbors which visually appear in the sampled identity. For instance, the sampled man in the first row shares a similar jawline to the nearest neighbor identity. The sampled identities also demonstrate the same ability as the original training identities to be composed into novel contexts. A variety of prompts are used in Fig. 12, yet the identities are consistent.
@@
Appendix B Composing Edits
We display additional examples of applying edits in w2w space based on the directions discovered using linear classifiers and CelebA labels. In Fig. 13, we demonstrate how the strength of these edits can be modulated and combined with minimal interference. These edits are apparent even in more complex scenes beyond face images. Also, the edits do not degrade other present concepts, such as the dog near the man (top left example).
In Figs. 14 and 15, we demonstrate how multiple edits can be progressively added in a disentangled fashion with minimal degradation to the identity. Additionally, since we operate in a subspace of weight space, these edits persist with a consistent appearance across different generations. For instance, even the man exhibits the edits as a painting in Fig. 14.
@@ Appendix C Inversion
Qualitative Inversion Comparison. In Figs. 16 and 17, we present qualitative comparisons of w2w inversion against Dreambooth trained with multiple images and a single image. Although mult-image Dreambooth slightly outperforms w2w inversion in identity preservations, its samples tend to lack realism compared to w2w inversion. We hypothesize that this may be due to using generated images for prior preservation and training on synthetic identity images. Dreambooth trained on a single image either generates an artifacted version of the original image or random identities. Notice how inversion into w2w space is even able to capture key characteristics of the child although babies are nearly to completely absent in the identites based on CelebA used to fine-tune our dataset of models.
@@ Appendix E Identity Datasets
In Fig. 19, we present examples of synthetic identity datasets used to conduct our Dreambooth fine-tuning as discussed in Sec 4. Each dataset is a set of ten images generated with [61] conditioned on a single CelebA [33] images associated with binary attribute labels. Note that we only display a subset of images per identity in the figure. Creating these synthetic datasets reduces intra-dataset diversity and creates a more consistent appearance for each subject. For instance, the first two rows in the figure are the same identity, but look drastically different. So we instead treat them as different identities associated with a different set of images.
Figure 19: Fine-tuning on synthetic examples allows Dreambooth fine-tuning to distill a consistent identity. The left column shows a CelebA image used to condition generation of a set of identity-consistent images in the right column associated with that identity. The consistent appearance of the identity enables a more consistent identity encoding.
@@ Appendix F Principal Component Basis
Figure 23: Identity inversion results with varying number of Principal Components. We optimize the coefficients for the first 1000, 10000, and 20000 Principal Component coefficients. Each column indicates a fixed generation seed and prompt. Inversion with the first 10,000 components balances parameter efficiency, realism, and identity preservation, and without overfitting to the single image.
Appendix G Timestep Analysis
@@
Figure 24: Injecting the edited weights at a lower timestep $T$ better preserves context at the expense of edit strength and fidelity.
Appendix H Extending to Other Domains
We find that similar subspaces can be created for other visual concepts beyond human identites. For instance, we apply the weights2weights framework to models encoding different dog breeds. We generate images with Stable Diffusion based on each of the $120$ dog classes from ImageNet [9]. We then conduct Dreambooth finetuning on each set of images to create a dataset of $120$ dog-encoding models. To find edit directions, we use GPT-4 [2] to create labels for each dog breed (e.g., wavy hair or not) and then train linear classifiers on the model weights like Sec. 3.3. We present examples of editing the dog-encoding models in Fig. 25. These hint at the generality of weights2weights, although more visual concepts should be explored to make a concrete conclusion.
Figure 25: Applying weights2weights edits on dog-encoding models. This suggests that weights2weights space can be constructed on domains beyond human identities.

### v3

@@ 6 Limitations
As with any data-driven method, w2w space inherits the biases of the data used to discover it. For instance, co-occurring attributes in the identity-encoding models would cause linear classifier directions to entangle them (e.g. gender and facial hair). However, as we scale the number of models, spurious correlations will drop as evidenced by Fig. 10. These semantic directions are also limited by the labels present in CelebA. Additionally, the span of the w2w space is dictated by the models used to create it. Thus, w2w space can struggle to represent more complex identities as seen in Fig. 12. Inversion in these cases amounts to projecting onto the closest identity on the weights manifold. Despite this, our analysis on the size of the model dataset reveals that forming a space using a larger and more diverse set of identity-encoding models can mitigate this limitation.
@@ Appendix A Sampling
We present additional examples of models sampled from w2w space in Fig. 13. The sampled models encode a diverse array of identities which are not copied from the dataset of model weights, as seen by comparing them to the nearest neighbor models from the training set. However, there are attributes borrowed from the nearest neighbors which visually appear in the sampled identity. For instance, the sampled man in the first row shares a similar jawline to the nearest neighbor identity. The sampled identities also demonstrate the same ability as the original training identities to be composed into novel contexts. A variety of prompts are used in Fig. 13 and the identities are still consistent.
@@
Appendix B Model Editing in w2w Space
Qualitative Results. We display additional examples of applying edits in w2w space based on the directions discovered using linear classifiers and CelebA labels. In Fig. 14, we demonstrate how the strength of these edits can be modulated and combined with minimal interference. These edits are apparent even in more complex scenes beyond face images. Also, the edits do not degrade other present concepts, such as the dog near the man (top left example).
In Figs. 15 and 16, we demonstrate how multiple edits can be progressively added in a disentangled fashion with minimal degradation to the identity. Additionally, since we operate in a subspace of weight space, these edits persist with a consistent appearance across different generations. For instance, even the man exhibits the edits as a painting in Fig. 15.
@@ Appendix B Model Editing in w2w Space
User Study. We present a two-alternative forced choice (2AFC) user study to evaluate the quality of identity edits in Tab. 3. Twenty-five users were given ten sets of images. Each set contained a randomly sampled original image of an identity, and then an image of that identity edited for an attribute using Concept Sliders [17], w2w, and text prompting. Users were then asked to choose between alternate pairs based on three criteria: identity preservation, alignment with the desired edit, and disentanglement. Our results in Tab. 3 show that users have a strong preference towards w2w edits. User instructions and an example question from the study are provided in Figs. 1718.
Table 3: User study on identity editing.
Method Win Rate (%) $\uparrow$
Sliders $28.4$
w2w $\mathbf{71.6}$
Prompting 12.812.8{12.8}12.8
w2w $\mathbf{87.2}$
Figure 17: Instructions provided to users for the identity editing user study.
Figure 18: Example question from the identity editing user study.
@@ Appendix C Inversion
Qualitative Inversion Comparison. In Figs. 19 and 20, we present qualitative comparisons of w2w inversion against Dreambooth trained with multiple images and a single image. Although mult-image Dreambooth slightly outperforms w2w inversion in identity preservations, its samples tend to lack realism compared to w2w inversion. We hypothesize that this may be due to using generated images for prior preservation and training on synthetic identity images. Dreambooth trained on a single image either generates an artifacted version of the original image or random identities. Notice how inversion into w2w space is even able to capture key characteristics of the child although babies are nearly to completely absent in the identites based on CelebA used to fine-tune our dataset of models.
@@
Comparison against Single Image Personalization Methods. We compare w2w inversion to single shot personalization methods Celeb-Basis [73] and IP-Adapter FaceID [^4] https://huggingface.co/h94/IP-Adapter-FaceID [72], following the same evaluation protocol from Sec. 4.4. These quantitative results are presented in Tab. 6. While Celeb-Basis optimizes in the input text embedding space for a given face image, IP-Adapter trains light-weight adapters [75, 39] to condition generation on any input face image.
We further conducted a two-alternative forced choice (2AFC) user study on the perceptual quality of w2w identity preservation. Twenty users were given ten sets of images. Each set contained a randomly sampled original image of the identity, and then three random images generated using IP-Adapter FaceID, Celeb-Basis, and w2w with the same random prompt. Users were then asked to choose between alternate pairs based on three criteria: identity preservation, prompt alignment, and diversity of generated images. Our results in Tab. 6 show that users found generations from w2w models capture identity better while also generating more diverse images that better align with the prompts.
Across both these metrics, w2w performs stronger than Celeb-Basis and IP-Adapter FaceID. Our results indicate that operating in our weight subspace is highly expressive and flexible as it is able to faithfully capture nuanced identity without overfitting to the input image. For instance, in Fig. 21, w2w inversion enables diverse generation with various poses, facial expressions, and clothing while maintaining identity.
Table 5: Single-shot personalization comparison.
Method ID Score $\uparrow$
Celeb-Basis ${0.60}\pm 0.02$
IP-Adapter FaceID $0.62\pm 0.02$
w2w $\mathbf{0.64}\pm 0.01$
Table 6: User study on identity inversion.
Method Win Rate (%) $\uparrow$
Celeb-Basis $13.4$
w2w $\mathbf{86.6}$
IP-Adapter FaceID 22.822.8{22.8}22.8
w2w $\mathbf{77.2}$
Figure 21: Qualitative comparison of single-shot personalization methods.
Below in Figs. 22 and 23, we provide the instructions provided to users for the identity inversion user study in addition to an example question.
Figure 22: Instructions provided to users for the identity inversion user study.
Figure 23: Example question from the identity inversion user study.
@@ Appendix E Identity Datasets
In Fig. 25, we present examples of synthetic identity datasets used to conduct our Dreambooth fine-tuning as discussed in Sec 4. Each dataset is a set of ten images generated with [67] conditioned on single CelebA [36] images associated with binary attribute labels. Note that we only display a subset of images per identity in the figure. The same identity can occur multiple times in different images in CelebA but have different appearances. So, creating these synthetic datasets reduces intra-dataset diversity and creates a more consistent appearance for each subject. For instance, the first two rows in the figure are the same identity, but look drastically different. So we instead treat them as different identities associated with a different set of images.
For evaluating identity edits from Sec. 4.3, we hold out $100$ identities, which results in leaving out ${\sim}{1000}$ models since multiple models may encode different instances of the same identity. For instance, if we left out the model encoding the identity in the first row of Fig. 25 for evaluation, the model encoding the second row identity would also be left out since it encodes the same identity but a different instance.
Figure 25: Fine-tuning on synthetic examples allows Dreambooth fine-tuning to distill a consistent identity. The left column shows a CelebA image used to condition generation of a set of identity-consistent images in the right column associated with that identity using [67]. The consistent appearance of the identity enables a more consistent identity encoding.
@@ Appendix F Principal Component Basis
Figure 29: Identity inversion results with varying number of principal components. We optimize the coefficients for the first $1000$, $10,000$, and $20,000$ Principal Component. Each column indicates a fixed generation seed and prompt. Inversion with the first $10,000$ components balances parameter efficiency, realism, and identity preservation without overfitting to the single image.
Figure 30: Identity preservation vs. number of principal components used for w2w inversion. We optimize the coefficients for the first $N$ principal components up to 20,000 and measure the average ID score for 100 inverted FFHQ identities.
Visualizing Principal Components. We provide a visualization of traversals along a set of principal components in Fig. 31. The principal components change attributes of the identity, although various semantic attributes are entangled. For instance, the first PC appears to change age, hair color, and hair style. The second PC appears to change gender and skin complexion. The third PC seems to change age, skin complexion, and facial hair. This motivates our use of linear classifiers to find separating hyperplanes in weight space and disentangle these attributes.
Figure 31: Traversal along the first three principal components in w2w space. The directions encode various entangled identity attributes such as age, gender, and facial hair.
Appendix G weights2weights for Other Visual Concepts
We find that similar subspaces can be created for other visual concepts beyond human identites. For instance, we apply the weights2weights framework to create two subspaces for models encoding different dog breeds and models encoding car types. We present examples of editing these models in Fig. 32. This suggest the generality of weights2weights and linear subspaces within diffusion model weights.
Figure 32: Applying weights2weights edits on dog-encoding models and car-encoding models. We create two datasets of model weights and apply PCA to define two separate weight subspaces. We then train linear classifiers to find semantic edit directions.
Appendix H Multi-Concept Merging
Multiple models living in w2w space cannot be merged since they live in the same weight subspace. So, merging will lead to interpolation of the identities. However, w2w models can be merged with models lying approximately orthogonal to the subspace. This merging can be done adding the weights. We present an example of merging a model living in w2w space with a model fine-tuned to encode “Pixar" style in Fig. 33.
Figure 33: Merging w2w models with non-identity models. Here, another model is fine-tuned to map $V_{2}$ to “Pixar” style. The two models are merged with simple addition.
Appendix I Timestep Analysis
@@
Figure 34: Injecting edited weights at varying timesteps. Using the edited weights at a smaller timestep $T$ better preserves context at the expense of edit strength and fidelity.